OpenStudy (insa):

A shop owner blends three types of tea. A,B and C, in the ratio 6:4:7. If the mass of tea C in the mixture is 28kg, find the difference in the masses of the other two types of tea.

OpenStudy (anonymous):

A would be 6/7 x 28kg = 24 kg B would be 4/7 x 28kg = 16 kg so the difference of the masses is 24 - 16 = 8kg